Article Title Performance Studies of the Monsoon Dataflow Processor

Authors James Hicks (Motorola Cambridge Research Center), Derek Chiou (Laboratory for Computer Science, MIT), Boon Seong Ang (Laboratory for Computer Science, MIT), and Arvind (Laboratory for Computer Science, MIT).

Performance Studies of the Monsoon Dataflow Processor

Author to send proofs to: Name: Derek Chiou mailing address: 545 Technology Square, Rm 203. email address: derek@abp.lcs.mit.edu Phone number: (617) 253-6871

In this paper, we examine the performance of Id, an implicitly parallel language, on Monsoon, an experimental dataflow machine. We briefly discuss the Monsoon architecture and our implementation of Id on Monsoon. Our run-time system and compiled Id code are general enough to run on any number of Monsoon processors without any change.

Our experiments show that speedups of more than seven are easily achieved on eight processors for most of the applications that we studied. We explain the sources of overhead that limit the speedup of each of our benchmark programs.

We also compare the performance of a single Monsoon processor with a DEC Station 5000 (MIPS R3000 processor), to understand the efficiency of Id execution on Monsoon. We find that the execution of Id programs on one Monsoon processor takes up to 3 times as many cycles as the corresponding C or Fortran programs executing on a MIPS R3000 processor. The statistics for the MIPS R3000 were obtained using pixie and excludes page fault and cache miss cycles. We identify the sources of inefficiency on Monsoon and suggest improvements, where possible. In many cases, however, improving single processor performance reduces parallel processor performance.